OpenStudy (anonymous):

Bob is moving and all of his sports cards are mixed up in a box. Twelve cards are baseball, eight are football, and five are basketball. If he reaches in the box and selects them at random, find each probability. 1. P(3 football) 2. P(3 baseball) 1) 8C3/25C3=56/2300 1) exactly the same 1) ?

OpenStudy (anonymous):

my answer 1 and 2 ?

OpenStudy (zarkon):

how many cards are being drawn...is it without replacement?

OpenStudy (anonymous):

It's an 8/25 for the first football card..

OpenStudy (anonymous):

25 total, drawn 3

OpenStudy (zarkon):

hmmm....;)

OpenStudy (zarkon):

if this is without replacement then the probability follows a hypergeometric distribution

OpenStudy (anonymous):

(8/25)(7/24)(6/23) right?

OpenStudy (zarkon):

yes ... which is the same as 8C3/25C3

OpenStudy (anonymous):

what do you mean 8C3/25C3

OpenStudy (anonymous):

8C3/25C3 yes correct ,that I did on frist one , secont one ?

OpenStudy (zarkon):

how many baseball cards are there ;)

OpenStudy (anonymous):

lol grrr.

OpenStudy (anonymous):

you mean football cards!? lol

OpenStudy (anonymous):

5

OpenStudy (zarkon):

2. P(3 baseball) "Twelve cards are baseball"

OpenStudy (anonymous):

12C3/25C3 ?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

I need more help Please! P(1 basketball, 2 football)

OpenStudy (zarkon):

ok...what is the denominator?

OpenStudy (anonymous):

25

OpenStudy (zarkon):

just 25?

OpenStudy (zarkon):

what was it in the other two problem when you took 3 cards?

OpenStudy (anonymous):

took 3 than 2 left

OpenStudy (zarkon):

no...the overall experiment is that you are picking 3 cards...how many ways can you do this?

OpenStudy (anonymous):

3!

OpenStudy (zarkon):

how many ways can I choose 3 cards from 25

OpenStudy (anonymous):

25!/(25-3)!3!

OpenStudy (zarkon):

yes...which is 25C3

OpenStudy (zarkon):

that is our denominator

OpenStudy (zarkon):

P(1 basketball, 2 football) lets break this up into two parts

OpenStudy (zarkon):

how many ways can we choose 1 basketball card from the 5 we have?

OpenStudy (anonymous):

5!/(5-1)!1!

OpenStudy (zarkon):

which is 5C1 good now how many ways can we choose 2 football cards from the 8 we have?

OpenStudy (anonymous):

8!/(8-2)!2!

OpenStudy (zarkon):

yes 8C2 multiply those two numbers we found and that is your numerator

OpenStudy (zarkon):

you get it?

OpenStudy (zarkon):

\[\frac{_5C_1\cdot _8C_2}{_{25}C_3}\]

OpenStudy (anonymous):

yes,how about : P(2 basketball, 1 baseball) 2 basketball = 5C2 1 baseball = 12C1

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

what denominator?

OpenStudy (zarkon):

\[\frac{_5C_2\cdot _{12}C_1}{_{25}C_3}\]

OpenStudy (zarkon):

the same

OpenStudy (zarkon):

you are performing the same experiment over and over....taking 3 cards

OpenStudy (anonymous):

I don't get is 25C3, I under stand 25 is total , 3?. How you get 3

OpenStudy (zarkon):

you are taking 3 cards

OpenStudy (anonymous):

3 card from the frist one ?

OpenStudy (zarkon):

prob=(number of ways to get a particular result from an experiment)/(number of ways to do the experiment)

OpenStudy (zarkon):

you wanted 2 basketball, 1 baseball that is 3 cards

OpenStudy (anonymous):

P(2 basketball, 1 baseball) deominator is 25C3 P(1 football, 2 baseball) deominator is 25C3 P(1 basketball, 1 football, 1 baseball) deominator is 25C3 P(2 baseball, 2 basketball) deominator is 25C4 P(2 football, 1 hockey) deominator is 25C3

OpenStudy (zarkon):

if that is all the info you are given then yes...those are correct

OpenStudy (anonymous):

thank, great help